how many ounces of 15% salt solution must be added to the 30 ounce of 70% salt solution to produce a 45% salt solution?
Answer by checkley77(12844)

.30*70+.15x=.45(30+x)
21+.15x=13.5+.45x
.15x-.45x=13.5-21
-.3x=-7.5
x=-7.5/-.3
x=25 oz. of 15% salt is used.
Proof:
.30*70+.15*25=.45(30+25)
21+3.75=.45*55
24.75=24.75